OpenStudy (anonymous):

if a,b,c r such that a+b+c=2 a^2+b^2+c^2=6 a^3+b^3+c^3=8 then what is a^4+b^4+c^4 equal to?

OpenStudy (anonymous):

Is it 9

OpenStudy (anonymous):

no

OpenStudy (anonymous):

OK D:

OpenStudy (anonymous):

b =2 , c = 1, a = -1?

OpenStudy (anonymous):

16 + 1 + 1 = 18

OpenStudy (anonymous):

u r right @Ishaan94

OpenStudy (anonymous):

Yay! but i would like to see algebraic solution.

OpenStudy (anonymous):

so that was hit and trial?

OpenStudy (anonymous):

Of course lol

OpenStudy (anonymous):

You get the ab +ac+bc = 1 from an expansion of (a+b+c)^2

OpenStudy (anonymous):

lolol i also solved it that way i too m luking fr algebraic sol

OpenStudy (anonymous):

@estudier , infact -1

OpenStudy (anonymous):

The next expansion gives you zeros

OpenStudy (anonymous):

@estudier just tell me how to use a^+b^2+c^2 the rest i can do myself

OpenStudy (anonymous):

And I am too lazy to rearrange the expansion of ^4 into zeros, -1's and so:-)

OpenStudy (anonymous):

i mean how to use the square terms to get a^4+b^4+c^4

OpenStudy (anonymous):

Has anybody managed to find abc value? If yes, then the problem is solved algebraically

OpenStudy (anonymous):

a*b*c = ?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

abc=-2

OpenStudy (anonymous):

Ok. Then I am getting the answer algebraically. Typing below

OpenStudy (anonymous):

@shivam_bhalla no need to type the whole sol.just tell how did u use a^+b^+c^ to get a^4+b^4+c^4

OpenStudy (anonymous):

(a+b+c)^4 = (a^2 + b^2 + c^2 -2)(a^2 + b^2 + c^2 -2) = (a^2+b^2+c^2)^2 + 4 - 4(a^2+b^2+c^2) = (a^4+b^4+c^4 + 2a^2b^2+ 2a^2c^2 + 2c^2b^2) + 4- 4*6 16 = a^4 + b^4 + c^4 + 2(a^2b^2+ a^2c^2 + c^2b^2) + 4 - 24

OpenStudy (anonymous):

(a+b+c)^2 = a^2 + b^2 + c^2 + 2ab+2bc+2ca (2)^2 = 6 + 2 (ab+bc+ca) ab+bc+ca = -1 (a^2+b^2+c^2)^2 = a^4+b^4+c^4+2a^b^2+2b^2c^2+2c^2a^2 6^2 = a^4+b^4+c^4 + 2 ((ab+bc+ca)^2 - 2ab^2c-2abc^2-2a^2bc) 36 = a^4+b^4+c^4 + 2 ((-1)^2 - 2abc(b+c+a)) 36 = a^4+b^4+c^4 + 2 (1 - 2(-2)(2)) a^4+b^4+c^4 = 18

OpenStudy (anonymous):

What did FoolForMath mean by Newton's Identity?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

thanks a lot @shivam_bhalla

OpenStudy (anonymous):

Welcome :D